2/3x-1/2x+1/3=3

The problem is suppose to read 2 over 3 x square this problem is ins fraction form. I dod not know how to put the exponent on to make it look square.

Answer by checkley75(3666)

(2/3)X^2-X/2+1/3=3
2X^2/3-X/2+1/3-3=0
2X^2/3-X/2-8/3=0
USING THE QUADRATIC EQUATION WE GET:
X=(1/2+-SQRT[-1/2^2-4*2/3*-8/3])/2*2/3
X=(1/2+-SQRT[1/4+64/9])/(4/3)
X=(1/2+-SQRT[7.36])/(4/3)
X=(1/2+-2.7)/(4/3)
X=(.5+2.7)/(4/3)
X=3.2(3/4)
X=2.4 ANSWER.
X=(.5-2.7)/(4/3)
X=-2.2(3/4)
X=-1.65 ANSWER.